Compute
1 - 2 + 3 - 4 + ... + 2005 - 2006
Thanks!

Answer by Aldorozos(172)

1-2=-1
3-4=-1
5-6=-1
.
.
2005-20056=-1
Here we have (2006/2)*(-1)= -1003